OpenStudy (anonymous):

8th grade math help!!!!!!!!! i will give metals!

OpenStudy (anonymous):

im in 7th but ill try :3

OpenStudy (anonymous):

Consider the following pair of equations: y = –2x + 8 y = x – 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form. (5 points)

OpenStudy (anonymous):

Consider the following pair of equations: –x – y = –5 y = x + 1 If the two equations are graphed, at what point do the lines representing the two equations intersect? (4 points) (3, 2) (3, –2) (–2, 3) (2, 3)

OpenStudy (anonymous):

Consider the following set of equations: Equation R: –3y = –3x - 9 Equation S: y = x + 3 Which of the following best describes the solution to the given set of equations? (4 points) No solution Infinite solutions Two solutions One solution

OpenStudy (anonymous):

Consider the following system of equations: y = 5x + 6 y = –x – 7 Which description best describes the solution to the system of equations? (4 points) Lines y = 5x + 6 and y = –x – 7 intersect the x-axis. Lines y = 5x + 6 and y = –x – 7 intersect the y-axis. Line y = 5x + 6 intersects line y = –x – 7. Line y = 5x + 6 intersects the origin.

OpenStudy (anonymous):

oh gosh i cant xD

OpenStudy (anonymous):

Two lines, A and B, are represented by the following equations: Line A: 2x + y = 6 Line B: x + y = 4 Which statement is true about the solution to the set of equations? (4 points) There is no solution. It is (4, 0). It is (2, 2). There are infinitely many solutions.

OpenStudy (anonymous):

Consider the following set of equations: Equation A: y = –x + 5 Equation B: y = 6x – 2 Which of the following is a step that can be used to find the solution to the set of equations? (4 points) –x + 5 = 6x - 2 –x = 6x + 2 –x + 5 = 5x –x – 2 = 6x + 5

OpenStudy (anonymous):

Jesus christ, calm down, 1 question at a time please.

OpenStudy (holly00d1248):

Aight you got to solve one of the equations for X or Y and it doesnt matter than plug that in and solve

OpenStudy (anonymous):

i have 6

OpenStudy (holly00d1248):

that confuse you or did i make sense

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

@holly00d1248